225 mL of a 1.45 mol/L solution of sodium iodide is added to 175 mL of a 0.85 mol/L solution of lead(II) nitrate

a) write the balanced chemical equation for this reaction

b) how many grams of solid precipitate will form in the process

c) what is the percent yield if 57.8g of precipitate was collected?


1
Expert's answer
2021-06-23T05:17:24-0400

a) 2NaI + Pb(NO3)2 = PbI2 + 2NaNO3


b) CM = n / V

n = CM x V

n (NaI) = 1.45 x 0.225 = 0.33 mol

n (Pb(NO3)2) = 0.85 x 0.175 = 0.15 mol

From the equation, n (NaI) = 2 x n (Pb(NO3)2).

Therefore, in this reaction Pb(NO3)2 is the limiting reactant.

n (PbI2) = n (Pb(NO3)2) = 0.15 mol

M (PbI2) = 461.01 g/mol

n = m / M

m (PbI2) = n x M = 0.15 x 461.01 = 69.15 g


c) w (PbI2) = 57.8 / 69.15 x 100 = 83.59%